OpenStudy (anonymous):

Find all solutions for the system of equations: (x^2)+2(y^2)=2 2(x^2)-3y=15

OpenStudy (experimentx):

x^2 = (15+3y)/2 (15+3y)/2 + 2y^2 = 2 4y^2+3y+11 = 0 ...i don't think you will have real solution

OpenStudy (anonymous):

i tried it, and when i graphed both, i go a parabola and a small curve, assuming that its a part of a circle, but they do not intersect

OpenStudy (experimentx):

an ellipse and a parabola

OpenStudy (anonymous):

for the first equation i got to y =sqrt((-x+2)/2) and second i got y= (2(x^2)/3)-5

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=%28x%5E2%29%2B2%28y%5E2%29%3D2%2C+2%28x%5E2%29-3y%3D15 they do not intersect ... you will have only imaginary solutions